OpenStudy (anonymous):

How do I do a substutition on this equation 7x+8y=0 2x+3y=-5 been working on it for 2hr need help please!

OpenStudy (anonymous):

express x in terms of y in equation (1) and plug it in equation (2) to find y then plug back to equation (1) to find x

OpenStudy (anonymous):

x=(-8y/7) -------eqn(1) 2x+3y =-5-------eqn(2) 2(-8y/7)+3y=-5 y= -7 then plu y into eqn (1) x=(-8*-7)/7 x=8 x=8, y=-7

OpenStudy (anonymous):

give me a medal

OpenStudy (anonymous):

@krypton will you help me please

OpenStudy (anonymous):

ok what is your question?

OpenStudy (anonymous):

i will give u a medal promise

OpenStudy (anonymous):

where is your question?

OpenStudy (anonymous):

@krypton; how did you get y=-7

OpenStudy (anonymous):

i hate when people just do it and ask for the medal with out fully explaining :/ 2(-8y/7)+3y=-5

OpenStudy (anonymous):

them after that what fo you do? im trying to figure how she got y=-7

OpenStudy (anonymous):

2(-8y/7)=-5-3y -16y/7=-5-3y -16y = -35 - 21y 5y = -35 y = -7

OpenStudy (anonymous):

understand? ^_^

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

@HawkCrimson what was not clear in my solution? i eve explained the steps before solving it for her uugh!!

OpenStudy (anonymous):

normally i explain stuffs,i dont solve or give answers.but decided to do it in her on case

OpenStudy (anonymous):

oh ok lol it was just the way right after u gave the explaination u said "give me a medal" lol

OpenStudy (anonymous):

it sounded like that's all u wanted from this

OpenStudy (anonymous):

well for me I have to be tought step by step on how to dew it see I had a whole packet but now im finally done with it cause of the help I got with the two equations I only needed I got them don't other equations on my packet done by myself juz needed help with two other equations and now im done I thank both of you